Accelerator Update for December 15, 2008 - December 17, 2008
Monday Dec 15
The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (13.3E10), with the Antiproton Source stacking (42.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store, and with the MiniBooNE and NuMI experiments taking beam.
Operations established store 6641 at 8:29 AM with an initial luminosity of 326.82E30.
At 9:55 AM, a Booster expert turned off a kicker (MKS07) to work on a timing problem. Operators adjusted the other kickers to make up for the missing kicker.
Operations reported at 3:57 PM that the Recycler had stashed antiprotons six times during the day shift.
At 5:30 PM, KRF3 tripped off. It reset.
Operations disabled the MTest beam at 6:10 PM.
Operations terminated store 6641 at 8:02 PM.
Operations prepared D0 and CDF for access at 8:20 PM.
D0 and CDF came out of access around 9:30 PM.
Operations put the TeV through a dry squeeze at 11:01 PM.
Operations put the TeV through a wet squeeze around 11:59 PM.
Operations reported at midnight that the Recycler had stashed nine times during the evening shift.
Tuesday Dec 16
The midnight shift began with the TeV in a wet squeeze, with Operations monitoring the stash (331.2E10), with Pbar stacking (26.5E10), and with NuMI and MiniBooNE taking beam.
Operations put the TeV through a dry squeeze at 12:24 AM.
Operations put the TeV into shot setup at 12:58 AM.
At 2:04 AM, the MI-60 lower power supply tripped off on an over current indication. Operations called in a power supply expert.
Operations put the TeV into a dry squeeze at 2:18 AM.
The MI power supply expert reported at 3:31 AM that there wasn’t an overcurrent problem.
Operations put the TeV back into shot setup at 4:22 AM.
Operations established store 6646 at 6:08 AM with an initial luminosity of 302.3E30.
Operations established beam to MTest at 6:29 AM.
Operations reported at 8:12 AM that the Recycler had stashed two times during the midnight shift.
A power supply expert reported at 10:02 AM that he had brought the MI-60 lower power supply back online.
Booster experts reported at 10:23 AM that they had turned off kickers MKS07 and 08 for repairs.
MI experts reported at 10:33 AM that they had begun repairs on MIRF2 and 17.
The Booster and MI accelerators resumed sending beam by 11:09 AM.
A MI expert began a study at 1:55 PM.
At 2:30 PM, Operations reported that the MI study was causing the MiniBooNE and MI beam permits to trip off.
The MI study ended at 3:22 PM.
Operations reported at 3:53 PM that the Recycler had stashed six times during the day shift.
At 6:02 PM, BRF19 tripped off. Operations called in an expert.
Operations put Pbar into standby at 6:32 PM.
At 7:19 PM, Booster RF experts reported that they had bypassed BRF19. Booster resumed sending beam.
Operations terminated store 6646 at 8:04 PM.
Operations disabled the MTest beam at 8:07 PM. The T988 experiment ended their run for this year. They will return for another run sometime next year.
Operations put the TeV into shot setup at 8:15 PM.
Operations established store 6647 at 9:54 PM with an initial luminosity of 296.36E30.
Operations reported at 11:58 PM that the Recycler had stashed five times during the evening shift.
Wednesday Dec 17
The midnight shift began with Operations monitoring store 6647 (192.83E30) and the stash (81.0E10), with Pbar stacking (33.8E10), and with NuMI and MiniBooNE taking beam.
At 3:26 AM, MIRF5 tripped off and caused the NuMI beam permit to also trip. Everything reset.
An interlock expert reported at 6:55 AM that he found nothing wrong with the ground fault monitor. It was a “soft” fault and the system had reset itself.
Operations began preparing the TeV for access at 7:18 AM. Vacuum technicians will work at sectors A47 and D47. Cryo system techs will check the sector D1 leads. TeV experts will work on the crystal collimator.
Operations reported at 8:09 AM that the Recycler had stashed eight times during the midnight shift.
The Plans for Wednesday
After the TeV access finishes, the plans for today are to resume stacking, stashing, and storing. The Run Coordinator will allow TeV, Pbar, and Recycler experts to conduct studies.
Comments and Suggestions
|last modified 01/05/2009 email Fermilab|
|Security, Privacy, Legal|